# Versatile Framework for Song Generation with Prompt-based Control

### v3

@@ Abstract
Song generation focuses on producing controllable high-quality songs based on various prompts. However, existing methods struggle to generate vocals and accompaniments with prompt-based control and proper alignment. Additionally, they fall short in supporting various tasks. To address these challenges, we introduce VersBand, a multi-task song generation framework for synthesizing high-quality, aligned songs with prompt-based control. VersBand comprises these primary models: 1) VocalBand, a decoupled model, leverages the flow-matching method for generating singing styles, pitches, and mel-spectrograms, allowing fast, high-quality vocal generation with style control. 2) AccompBand, a flow-based transformer model, incorporates the Band-MOE, selecting suitable experts for enhanced quality, alignment, and control. This model allows for generating controllable, high-quality accompaniments aligned with vocals. 3) Two generation models, LyricBand for lyrics and MelodyBand for melodies, contribute to the comprehensive multi-task song generation system, allowing for extensive control based on multiple prompts. Experimental results demonstrate that VersBand performs better over baseline models across multiple song generation tasks using objective and subjective metrics. Audio samples are available at https://aaronz345.github.io/VersBandDemo.
@@ 1 Introduction
Song generation focuses on producing complete musical pieces based on text prompts (about lyrics, melodies, singing, and music styles) and optional audio prompts. Unlike singing voice synthesis (SVS) (Zhang et al., 2024b; Guo et al., 2025), which focuses on the singing component, or music generation (Dong et al., 2018; Agostinelli et al., 2023) for synthesizing only instrumental tracks, song generation involves synthesizing both high-quality vocals and accompaniments with high-level prompt-based control and proper alignment in rhythm, melody, and beats. (Li et al., 2024a).
@@
To address these challenges, we introduce VersBand, a multi-task song generation framework for synthesizing high-quality, aligned songs with prompt-based control. Following the human perception that accompaniment complements vocal melody with complex harmonic and rhythmic structure (Zhiqing et al., 2024), we generate them separately. To achieve fast and high-quality vocal generation with control, we design a decoupled model VocalBand, predicting singing styles, pitches, and mel-spectrograms based on the flow-matching method. Based on the complex nature of music, we introduce a flow-based transformer model AccompBand to generate high-fidelity, controllable, aligned accompaniments. We design Band-MOE (Mixture of Experts), selecting suitable experts for enhanced quality, alignment, and control. Additionally, we add two generation models, LyricBand for lyric generation and MelodyBand for melody generation, contributing to the comprehensive multi-task song generation system. Our experiments on both open-source and web-crawled Chinese and English song datasets show that VersBand can generate high-quality songs with control, outperforming other baseline models in multiple related song generation tasks. The main contributions of our work are summarized as follows:
@@ 2.1 Singing Voice Synthesis
Singing Voice Synthesis (SVS) rapidly advances for generating singing voices from given lyrics and music scores. Choi and Nam (2022) presents a melody-unsupervised model, eliminating the need for temporal alignment. VISinger 2 (Zhang et al., 2022b) employs digital signal processing techniques to enhance fidelity, while Kim et al. (2024) uses adversarial multi-task learning to improve the singing naturalness. StyleSinger (Zhang et al., 2024a) facilitates style transfer by extracting styles via a residual quantization method. Additionally, PromptSinger (Wang et al., 2024) attempts to control speaker identity based on text descriptions. GTSinger (Zhang et al., 2024c) releases a dataset with multiple style annotations. Despite these advancements, these approaches can not generate aligned accompaniment. Melodist (Zhiqing et al., 2024) has introduced a text-to-song model that sequentially generates vocals and accompaniments using auto-regressive transformers. TCSinger (Zhang et al., 2024b) and TechSinger (Guo et al., 2025) utilize input sequences to control vocal techniques. However, achieving high-quality vocal generation with high-level control by natural language text or audio prompts remains challenging.
@@ 2.2 Accompaniment Generation
Research in accompaniment usually focuses on musical symbolic tokens. MuseGAN (Dong et al., 2018) employs a GAN-based approach to generate symbolic music. SongMASS (Sheng et al., 2021) uses transformer models to generate lyrics or melodies conditioned on each other. MusicLM (Agostinelli et al., 2023) leverages joint textual-music representations from MuLan (Huang et al., 2022) to generate semantic and acoustic tokens using transformer decoders. MusicLDM (Chen et al., 2024) incorporates beat-tracking information to address potential plagiarism concerns in music generation. Additionally, SingSong (Donahue et al., 2023) introduces a model that generates background music to complement the provided vocals. SongCreator (Lei et al., 2024) can create songs based on lyrics and audio prompts. Recently, MelodyLM (Li et al., 2024a) has employed transformer and diffusion models for decoupled song generation. Nevertheless, challenges remain in generating high-quality music with effective control based on multiple prompts. Existing methods also lack mechanisms for alignment with vocals and support for multiple song generation tasks.
@@ 3 Method
This section introduces VersBand. We design two distinct models, VocalBand for vocals and AccompBand for accompaniments, tailored to their unique characteristics. Additionally, we include LyricBand for lyrics and MelodyBand for melodies, composing a multi-task song generation system.
@@ 3.1 Multi-Task Song Generation
As shown in Figure 2 (a), VersBand handles multi-task song generation based on text and audio prompts. We employ a text encoder to generate text tokens $z_{p}$. When lyrics or music scores are not provided, LyricBand and MelodyBand predict phonemes $p$ and notes $n$ (pitch and duration) as target contents. Next, in Figure 2 (b), to achieve fast and high-quality vocal generation with granular control, we introduce VocalBand, which decouples the content $z_{c}$, timbre $z_{t}$, and style $z_{s}$. Through the Flow-based Pitch Predictor, Mel Decoder, and pre-trained vocoder, the target vocal $y_{v}$ is synthesized. Then, in Figure 2 (c), for the complex nature of accompaniment, we design AccompBand to achieve superior quality, alignment, and control. AccompBand uses two encoders to extract embeddings $z_{v}$ from $y_{v}$ and $x$ from ground truth (GT) accompaniment $\hat{y_{a}}$ during training. $z_{v}$ and noise-injected $x_{t}$ are processed by Band Transformer Blocks with Band-MOE, which selects suitable experts by $z_{v}$, $z_{p}$, and time step $t$ for enhanced quality, alignment, and control. During inference, the ordinary differential equation (ODE) solver, accomp decoder, and vocoder generate the target accompaniment $y_{a}$. $y_{v}$ and $y_{a}$ are combined to the final target song $y$.
@@ 4.4 Song Generation
Melodist 3.81 1.34 0.39 84.12$\pm$1.54 85.97$\pm$1.51 74.86$\pm$1.13
MelodyLM 3.42 1.35 0.35 85.23$\pm$1.62 86.44$\pm$0.90 75.41$\pm$1.34
VersBand (w/o lyrics) 3.37 1.30 0.50 86.65$\pm$0.91 85.98$\pm$1.08 77.02$\pm$1.33
VersBand (w/o scores) 3.38 1.31 0.48 85.12$\pm$0.77 85.15$\pm$1.22 75.91$\pm$1.62
VersBand (w/o prompts) 3.55 1.35 - 83.49$\pm$1.20 - 74.87$\pm$1.68
VersBand (w/ full) 3.01 1.27 0.58 87.92$\pm$1.73 88.03$\pm$0.59 80.51$\pm$1.66
Table 4: Results of song generation. Content includes lyrics and music scores.
@@ 4.5 Ablation Study
We remove various components from text prompts for evaluation. As shown in Table 4, even with a minimum input, VersBand still delivers remarkable performance. When listening to songs generated for various tasks on our demo page, it is evident VersBand shows strong controllability and expressiveness over various text prompts, along with the ability to produce intricate, skillful vocals employing multiple techniques, and complex, well-aligned accompaniments featuring harmonious instrumentation. For more ablation studies, please refer to Appendix J.
@@ 5 Conclusions
In this paper, we present VersBand, a multi-task song generation framework for synthesizing high-quality, aligned songs with prompt-based control. We mainly design these models: 1) VocalBand, a decoupled model leveraging the flow-matching model for singing styles, pitches, and mel-spectrograms generation, allowing fast and high-quality vocal generation with style control. 2) AccompBand, a flow-based transformer model, incorporates the Band-MOE, selecting suitable experts for enhanced quality, alignment, and control. This model generates controllable, high-quality accompaniments aligned with vocals. 3) Two generation models, LyricBand for lyrics and MelodyBand for melodies, contribute to the comprehensive multi-task song generation system. Experimental results demonstrate that VersBand performs better over baseline models across multiple song generation tasks using objective and subjective metrics.
@@ 7 Ethics Statement
Large-scale generative models always present ethical challenges. VersBand, due to its multi-task song generation capabilities, could potentially be misused for dubbing in entertainment short videos, raising concerns about the infringement of famous singers’ copyrights. Then, its ability to transfer and control multiple song styles about lyric, melody, singing, and music, lowers the requirements for high-quality, personalized, controllable song generation, posing some risks like unfair competition and potential unemployment for professionals in related music and singing occupations. To mitigate these potential risks, we will explore methods like music watermarking to protect individual privacy.
@@ B.1 Model Details
Our VersBand framework consists of four models: VocalBand, AccompBand, LyricBand, and MelodyBand. For the text encoder, we use FLAN-T5-large (Chung et al., 2024), while we also test BERT-large (Devlin et al., 2018) and the text encoder of CLAP (Elizalde et al., 2023) in Appendix J.1. Our vocoder is the pre-trained HiFi-GAN (Kong et al., 2020). For detailed hyperparameters of each component, please refer to Appendix C.1, D.1, E.1, and E.2.
@@ I.4 Vocal-to-Song Generation
Methods FAD $\downarrow$ KLD $\downarrow$ CLAP $\uparrow$ OVL $\uparrow$ REL $\uparrow$
VersBand (T5) 3.01 1.27 0.58 87.92$\pm$1.73 88.03$\pm$0.59
VersBand (CLAP) 3.31 1.34 0.41 85.36$\pm$1.57 86.03$\pm$1.39
VersBand (BERT) 3.12 1.29 0.49 87.02$\pm$0.84 87.21$\pm$0.83
Table 13: Results of ablation study on different text encoders.
Figure 12: The statistics of global routing and acoustic routing in Band-MOE.
@@ I.5 Accompaniment-to-Song Generation
Appendix J Ablation Study
J.1 Experiments on Text Encoder
For the text encoder, following previous work (Zhiqing et al., 2024), we test FLAN-T5-large (Chung et al., 2024), BERT-large (Devlin et al., 2018), and the text encoder of CLAP (Elizalde et al., 2023). Table 13 shows that we test VersBand without inputting lyrics or music scores. It can be seen that the T5 text encoder outperforms the other two text encoders in both quality and relevance, but has only a slight advantage over BERT, which is likely due to its larger parameter count and multi-task capability.
J.2 Experiments on MOE
To demonstrate the effectiveness of our MOE, we conducted experiments on the final routing behavior. As shown in Figure 12 (a), we can observe that our global routing behaves as expected. As the noise level decreases, the weighting of outputs from Aligned MOE and Controlled MOE changes accordingly: 1) At early time steps (near 0), where the hidden representation $h$ is highly noisy, the network prioritizes matching with the vocal for coherent reconstruction, thus the weight of the Aligned MOE is higher. 2) At later time steps (near 1), where $h$ has been largely reconstructed, the network focuses more on refining stylistic details, relying heavily on text prompts, thus the weight of the Controlled MOE is higher.
As shown in Figure 12 (b), the Acoustic MOE also behaves as expected by assigning different experts to different channels. We encode the mel-spectrogram into 20 dimensions through the accomp encoder, resulting in 20 channels and selecting experts for each channel. We perform a statistical analysis of the softmax output probabilities before expert selection. 1) Expert 1 focuses on channels 0 to 7, which include instruments that provide foundational rhythm and depth, such as bass guitars, kick drums, low-frequency percussion, and the lower registers of piano and organ. 2) Expert 2 specializes in channels 4 to 12, capturing the richness of rhythm guitars, mid-range piano notes, and various percussion instruments that contribute to the fullness and body of the music. 3) Expert 3 targets channels 9 to 16, encompassing lead guitars, higher piano octaves, string instruments, and brass instruments. This allows the model to capture melodic elements and intricate harmonics that enhance the expressiveness of the accompaniment. 4) Expert 4 is assigned to channels 14 to 19, focusing on cymbals, hi-hats, flutes, and high-frequency string overtones that contribute to the brightness and airiness of the music.
Appendix K MIDI and F0
The pitch predicted by VocalBand refers to the F0 of the vocal, while MelodyBand predicts the MIDI music score. These two representations are fundamentally different in terms of both their structure and level of granularity.
F0 (Fundamental Frequency)
refers to the lowest frequency of a periodic waveform, typically representing the pitch of a sound. In vocal music, it corresponds to the pitch produced by the singer’s voice. f0 is continuous and fine-grained, capturing subtle pitch variations that occur naturally during singing. This includes slight pitch bends, vibrato, and dynamic vocal expressions. The vocal f0 prediction in VocalBand focuses on capturing these nuanced and continuous pitch variations that define vocal performances.
MIDI (Musical Instrument Digital Interface)
is a digital representation of music that encodes musical notes, chords, velocities, and other musical events in a discrete, symbolic format. MIDI defines music in terms of note numbers (representing pitch) and timing information (e.g., note onset and duration). Unlike f0, MIDI is typically coarser in its resolution. It represents pitch as discrete notes (e.g., C4, D5), with fixed intervals, meaning it does not capture the smooth pitch fluctuations found in vocal performances or continuous instruments. The MIDI music score predicted by MelodyBand works with these discrete representations, focusing more on the overall structure of the music (i.e., sequence of notes and timing) rather than the fine details of pitch fluctuations. In summary, f0 is a detailed, continuous signal that reflects the pitch of vocal sounds, whereas MIDI is a more abstract, discrete representation of musical notes with a lower level of granularity. These differences illustrate the challenge and complexity of predicting both types of information, each serving a distinct purpose in music generation and synthesis.
Modeling vocal MIDI as an intermediate modality for supervision can significantly improve the synthesis performance of models. Directly predicting f0 in VocalBand would complicate the modeling of phoneme duration and the f0 itself, leading to poorer results and making it harder for the model to perform core synthesis tasks. Additionally, singing voice synthesis models typically require user input of music scores for coarse control. To maintain consistency within our model, we opted to include MIDI prediction.

### v5

@@ Abstract
Song generation focuses on producing controllable high-quality songs based on various prompts. However, existing methods struggle to generate vocals and accompaniments with prompt-based control and proper alignment. Additionally, they fall short in supporting various tasks. To address these challenges, we introduce VersBand, a multi-task song generation framework for synthesizing high-quality, aligned songs with prompt-based control. VersBand comprises these primary models: 1) VocalBand, a decoupled model, leverages the flow-matching method for generating singing styles, pitches, and mel-spectrograms, allowing fast, high-quality vocal generation with style control. 2) AccompBand, a flow-based transformer model, incorporates the Band-MOE, selecting suitable experts for enhanced quality, alignment, and control. This model allows for generating controllable, high-quality accompaniments aligned with vocals. 3) Two generation models, LyricBand for lyrics and MelodyBand for melodies, contribute to the comprehensive multi-task song generation system, allowing for extensive control based on multiple prompts. Experimental results show that VersBand outperforms baseline models across multiple song generation tasks using objective and subjective metrics. Demos and codes are available at https://aaronz345.github.io/VersBandDemo and https://github.com/AaronZ345/VersBand.
@@ 1 Introduction
Song generation focuses on producing complete musical pieces based on text prompts (about lyrics, melodies, singing, and music styles) and optional audio prompts. Unlike singing voice synthesis (SVS) (Zhang et al., 2024b), which focuses on the vocal component, or music generation (Dong et al., 2018) for synthesizing only instrumental tracks, song generation involves synthesizing both high-quality vocals and accompaniments with high-level prompt-based control and proper alignment in rhythm, melody, and beats Li et al. (2024a).
@@
To address these challenges, we introduce VersBand, a multi-task song generation framework for synthesizing high-quality, aligned songs with prompt-based control. Following the perception that accompaniment complements vocal melody with complex harmonic and rhythmic structure (Zhiqing et al., 2024), we generate them separately. To achieve fast and high-quality vocal generation with control, we design a decoupled model, VocalBand, predicting singing styles, pitches, and mel-spectrograms based on the flow-matching method. Based on the complex nature of music, we introduce a flow-based transformer model, AccompBand, to generate high-fidelity, controllable, aligned accompaniments. We design Band-MOE (Mixture of Experts), selecting suitable experts for enhanced quality, alignment, and control. Additionally, we add two generation models, LyricBand for lyric generation and MelodyBand for melody generation, contributing to the comprehensive multi-task song generation system. Our experiments show that VersBand can generate high-quality songs with control, outperforming other baseline models in multiple related song generation tasks. The main contributions are summarized as follows:
@@ 2.1 Singing Voice Synthesis
Singing Voice Synthesis (SVS) rapidly advances for generating singing voices from given lyrics and music scores. Choi and Nam (2022) presents a melody-unsupervised model, eliminating the need for temporal alignment. VISinger 2 (Zhang et al., 2022b) employs digital signal processing techniques to enhance fidelity, while Kim et al. (2024) uses adversarial multi-task learning to improve the singing naturalness. StyleSinger (Zhang et al., 2024a) facilitates style transfer by extracting styles via a residual quantization method. GTSinger (Zhang et al., 2024c) releases a dataset with multiple style annotations. Despite these advancements, these approaches can not generate aligned accompaniment. Melodist (Zhiqing et al., 2024) has introduced a text-to-song model that sequentially generates vocals and accompaniments using auto-regressive transformers. TCSinger (Zhang et al., 2024b) and TechSinger (Guo et al., 2025) utilize input sequences to control vocal techniques. However, achieving high-quality vocal generation with high-level control by natural language text or audio prompts remains a challenging task.
@@ 2.2 Accompaniment Generation
Research in accompaniment usually focuses on musical symbolic tokens. MuseGAN (Dong et al., 2018) employs a GAN-based approach to generate symbolic music. SongMASS (Sheng et al., 2021) uses transformer models to generate lyrics or melodies conditioned on each other. MusicLM (Agostinelli et al., 2023) leverages joint textual-music representations from MuLan (Huang et al., 2022) to generate semantic and acoustic tokens using transformer decoders. MusicLDM (Chen et al., 2024) incorporates beat-tracking information to address potential plagiarism concerns in music generation. SongCreator (Lei et al., 2024) can create songs based on lyrics and audio prompts. Recently, MelodyLM (Li et al., 2024a) has employed transformer and diffusion models for decoupled song generation. Nevertheless, challenges remain in generating high-quality music with effective control based on multiple prompts. Existing methods also lack mechanisms for alignment with vocals and support for multiple song generation tasks.
@@ 3.1 Multi-Task Song Generation
As shown in Figure 2 (a), VersBand handles multi-task song generation based on text and audio prompts. We design two distinct models, VocalBand for vocals and AccompBand for accompaniments, tailored to their unique characteristics. First, we employ a text encoder to generate text tokens $z_{p}$. When lyrics or music scores are not provided, LyricBand and MelodyBand predict phonemes $p$ and notes $n$ (pitch and duration) as target contents. Next, in Figure 2 (b), to achieve fast and high-quality vocal generation with granular control, we introduce VocalBand, which decouples the content $z_{c}$, timbre $z_{t}$, and style $z_{s}$. Through the Flow-based Pitch Predictor, Mel Decoder, and pre-trained vocoder, the target vocal $y_{v}$ is synthesized. Then, in Figure 2 (c), for the complex nature of accompaniment, we design AccompBand to achieve superior quality, alignment, and control. AccompBand uses two encoders to extract embeddings $z_{v}$ from $y_{v}$ and $x$ from ground truth (GT) accompaniment $\hat{y_{a}}$ during training. $z_{v}$ and noise-injected $x_{t}$ are processed by Band Transformer Blocks with Band-MOE, which selects suitable experts by $z_{v}$, $z_{p}$, and time step $t$ for enhanced quality, alignment, and control. During inference, the ordinary differential equation (ODE) solver, accomp decoder, and vocoder generate the target accompaniment $y_{a}$. $y_{v}$ and $y_{a}$ are combined to the final target song $y$.
@@ 4.4 Song Generation
[…]
Table 4: Results of song generation.
@@ 4.5 Ablation Study
We remove various components from text prompts for evaluation. As shown in Table 4, even with a minimum input, VersBand still delivers remarkable performance. When listening to songs generated for various tasks on our demo page, it is evident that VersBand shows strong controllability and expressiveness over various text prompts, along with the ability to produce intricate, skillful vocals employing multiple techniques, and complex, well-aligned accompaniments featuring harmonious instrumentation.
@@ 5 Conclusions
In this paper, we present VersBand, a multi-task song generation framework for synthesizing high-quality, aligned songs with prompt-based control. We mainly design these key models for the multi-task comprehensive song generation system: 1) VocalBand, a decoupled model leveraging the flow-matching model for singing styles, pitches, and mel-spectrograms generation, allowing fast and high-quality vocal generation with style control. 2) AccompBand, a flow-based transformer model, incorporates the Band-MOE, selecting suitable experts for enhanced quality, alignment, and control. This model generates controllable, high-quality accompaniments aligned with vocals. 3) Two generation models, LyricBand for lyrics and MelodyBand for melodies, contribute to the versatile song generation system. Our experimental results demonstrate that VersBand outperforms baseline models across multiple song generation tasks, as measured by both objective and subjective metrics.
@@ 7 Ethics Statement
Large-scale generative models always present ethical challenges. VersBand, due to its multi-task song generation capabilities, could potentially be misused for dubbing in entertainment short videos, raising concerns about the infringement of famous singers’ copyrights. Then, its ability to transfer and control multiple song styles about lyric, melody, singing, and music lowers the requirements for high-quality, personalized, controllable song generation, posing some risks like unfair competition and potential unemployment for professionals in related music and singing occupations. To mitigate these potential risks, we will explore methods like music watermarking to protect individual privacy.
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant No. 62222211, National Key R&D Program of China under Grant No.2022ZD0162000, and National Natural Science Foundation of China under Grant No.U24A20326.
@@ B.1 Model Details
Our VersBand framework consists of four models: VocalBand, AccompBand, LyricBand, and MelodyBand. For the text encoder, we use FLAN-T5-large (Chung et al., 2024). Our vocoder is the pre-trained HiFi-GAN (Kong et al., 2020). For detailed hyperparameters of each component, please refer to Appendix C.1, D.1, E.1, and E.2.